# Strong coupling between a single photon and a photon pair

### v2

@@ Abstract
Figure 4: |||| Interference between two probe tones. a(c), Gain of the transmitted amplitude of the signal field in the $n=1$ ($n=2$) mode of the resonator versus the average photon number ${\overline{n}}_{\mathrm{2(1)}}$ and the phase of the control field in the $n=2$ ($n=1$) mode of the resonator. The average photon numbers of the signal fields in the resonator’s $n=1$ and $n=2$ modes are about 0.25 and 0.13, respectively. The frequencies of the probe tones are $\omega/2\pi=4.905$ GHz and $2\omega/2\pi$ (at ${\delta\mathrm{\Phi}}_{\mathrm{ext}}=-46~{}{\rm m\Phi_{0}}$). b (d), Theoretical calculations corresponding to the plots in a (c). The phase shifts in the experimental data in a (c) are an artifact due to an electronic delay.
@@
To further characterize the nature of the strong nonlinear coupling, we measured the SHG at the single-photon power level by applying a weak signal at $\omega$ and monitoring the SHG signal at $2\omega$. The result is shown in Fig. 3. As expected, the SHG signal is more intense near the minimum of the avoided-level crossing and for flux offsets where the system is double resonant: $\tilde{\omega}_{3}=2\tilde{\omega}_{1}$, as confirmed (despite some shift due to some small discrepancy in the fitted energy levels) by the theoretical calculations in Fig. 3b. Figure 3c shows that the SHG amplitude is thresholdless and starts with a linear dependence on the mean value of input photons $\bar{n}_{1}$, while it tends to saturate for higher input powers. It can clearly be distinguished from the noise floor even for the mean number of input photons in the resonator ${\overline{n}}_{1}$ significantly below $1$, which is a striking unprecedented result, made possible via the photonic strong coupling regime shown in Fig. 2b. The mean number of input photons in the resonator is evaluated using a standard input–output relationship Milburn2008Springer (see Supplementary Information for details). The hybridization between one- and two-photon states at the minimum of the avoided-level crossing gives rise to matrix elements ${X}_{+,1},\,{X}_{+,0},\,{X}_{1,0}$ all of the order of unity, which enable observation of the SHG well below the single-photon power level.
@@
A further direct consequence of the achieved photonic strong coupling is the observation of interference between a probe at $\omega$ ($2\omega$) and a control tone at $2\omega$ ($\omega$) both at the single-photon power level, resulting in the amplification or suppression of the transmitted amplitude of probe signals. The amplification and suppression are phase dependent, just as in the case of degenerate parametric amplifiers Tsai2008APL ; Lehnert2008NP . The difference is that a much stronger pump tone is not needed here, and we observe these interference phenomena at an incredibly low-power limit (average photon number in the resonator below one). The results are shown in Fig. 4. Changing the phase of the control tones, we can see that the transmitted amplitude of the signal at the $n=2$ ($n=1$) mode (signal field) is amplified or suppressed periodically with visibility approaching 1. It turns out that the visibility of the interference for the signal at $2\omega$ becomes relevant for higher intensities of the control signal (at $\omega$) (see Fig. 4c,d) with respect to what is observed for the signal at $\omega$ (see Fig. 4a,b). Conversely, to obtain a high visibility for interference at $\omega$, it requires a larger probe signal at $\omega$. This difference originates from the fact that the interference at $2\omega$ requires the stimulated conversion of two photons of the control tone. On the contrary, the probe signal at $\omega$ receives photon pairs from the down conversion of photons at $2\omega$ of the control tone.
@@ Discussion
We experimentally observed an unstimulated strong coupling between a single photon and a photon pair. The observed single-photon strong nonlinear coupling is enabled by an embedded-in flux qubit, which, while remaining in its ground state, acts as an effective coupler between two modes of a superconducting coplanar waveguide resonator. The strong coupling between photon states of opposite parity is enabled by the spontaneous mixing of even- and odd-number photon states, which is a unique feature of the generalized quantum Rabi Hamiltonian in the ultrastrong coupling regime Nori2017PRA ; Wang2023NC .
The counterpart in time-resolved experiments of the quantum Rabi splitting reported here are quantum Rabi oscillations enabling the conversion of a single photon into a photon pair and vice versa with a period $T_{R}=2\pi/g_{\rm eff}$ and with an efficiency that can potentially approach $100\%$ Nori2017PRA (for $g_{\rm eff}$ significantly larger than the losses). In contrast to most conventional implementations of nonlinear optics, the regime demonstrated here can reach unit efficiency only using a minimal number of photons and without exciting atomic transitions, which introduce additional losses. For example, in many cases pure dephasing in superconducting artificial atoms is the main source of decoherence. These results unlock a new regime for photons in quantum nonlinear optics Lukin2014NPh ; Bouwmeester2006PRL (coherent interaction between photon multiplets), which has important implications in both fundamental research Nori2017PRA and applications Zeilinger2011Nature ; Niu2018PRL . Examples of direct applications include (i) the realization of deterministic quantum light (single photon and photon pair) sources as in the cases of traditional spontaneous parametric down conversion (SPDC) but with an unprecedented efficiency (approaching $100\%$); (ii) an all-optical deterministic controlled-phase gate ($|\widetilde{0,1}\rangle\leftrightarrow-|\widetilde{0,1}\rangle$, a geometric phase effect under a full Rabi oscillation ($2T_{R}$), where the presence or the absence of a photon in mode 1 corresponds to the control qubit). This two-photon quantum phase gate represents a key element for the realization of a still elusive universal nonlinear optical quantum computing (NOQC) Zeilinger2011Nature ; Niu2018PRL . These controlled quantum Rabi oscillations can be realized by using suitable flux-offset pulses to tune the system in and out the minimum of the avoided level crossing $|\widetilde{0,1}\rangle\leftrightarrow|\widetilde{2,0}\rangle$.
@@ Methods
Our device is a 3 mm $\times$ 10 mm chip fabricated on a single crystal sapphire substrate (see Supplementary Fig. 1b). The superconducting coplanar waveguide resonator is defined by patterning a niobium thin film using optical lithography. The flux qubit consists of two identical large JJs and a smaller JJ (with an area ratio $\alpha\approx\mathrm{0.5}$) in a loop geometry. The qubit loop and its connection to the center conductor of the coplanar waveguide resonator are made by electron beam lithography and double-angle evaporation of aluminum. Placing the qubit loop near the antinode of the current distribution for the $n=2$ mode of the resonator makes the qubit coupled to the $n=1$ mode and the $n=2$ mode of the resonator simultaneously (see Fig. 1a). The values of the Josephson energy $E_{J}/2\pi=173\ \mathrm{GHz}$, the area ratio $\alpha\approx\mathrm{0.5}$ and ${E_{J}}/{E_{C}}\approx 100$ of the JJs in the qubit are optimized to ensure the flux qubit working as an effective nonlinear coupler, i.e., both far-away detuned and strongly coupled with the relevant modes of the resonator, and are used to numerically diagonalize the three-junction flux qubit Hamiltonian to estimate the values of $\Delta$ and $I_{p}$.
@@
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
@@
The code that support the findings of this study are available from the corresponding authors upon reasonable request.
Author contributions
S.W., T.L. and J.Q.Y. conceived the experiment. Y.W. and M.C. fabricated the sample. S.W. performed the experiment. S.W., A. M., A.R., and S.S. analyzed and interpreted the data. A.M., A.R., F.N., and S.S. developed the theory. All authors discussed the results and contributed to the writing of the manuscript.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China (Grant No. 2022YFA1405200), the National Natural Science Foundation of China (Grants No. 92265202, No. 11934010, No. 92365210, No. 62074091, No. 62365210, No. U2230402, No. 12374325, and No. 12304387), and Tsinghua University Initiative Scientific Research Program. A.R. acknowledges the QuantERA grant SiUCs (Grant No. 731473), the PNRR MUR project PE0000023-NQSTI, and the ICSC - Centro Nazionale di Ricerca in High-Performance Computing, Big Data and Quantum Computing. Y.L. acknowledges the support of Beijing Municipal Science and Technology Commission (Grant No. Z221100002722011), and Young Elite Scientists Sponsorship Program by CAST (Grant No. 2023QNRC001). S.S. acknowledges the Army Research Office (ARO) (Grant No. W911NF1910065). F.N. is supported in part by Nippon Telegraph and Telephone Corporation (NTT) Research, the Japan Science and Technology Agency (JST) [via the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&D Grant No. JPMJMS2061], the Asian Office of Aerospace Research and Development (AOARD) (Grant No. FA2386-20-1-4069), and the Office of Naval Research (ONR).

### v3

@@ Results
Figure 4: $|$ Interference between two probe tones. a(c), Gain of the transmitted amplitude of the signal field in the $n=1$ ($n=2$) mode of the resonator versus the average photon number ${\overline{n}}_{\mathrm{2(1)}}$ and the phase of the control field in the $n=2$ ($n=1$) mode of the resonator. The average photon numbers of the signal fields in the resonator’s $n=1$ and $n=2$ modes are about 0.25 and 0.13, respectively. The frequencies of the probe tones are $\omega/2\pi=4.905$ GHz and $2\omega/2\pi$ (at ${\delta\mathrm{\Phi}}_{\mathrm{ext}}=-46~{\rm m\Phi_{0}}$). b (d), Theoretical calculations corresponding to the plots in a (c). The gain is defined as the ratio between the output signal amplitude at $\omega$ ($2\omega$) when a secondary signal is applied at $2\omega$ ($\omega$), and the output amplitude of the same signal at $\omega$ ($2\omega$) in the absence of the secondary signal applied at $2\omega$ ($\omega$). The phase shifts in the experimental data in a (c) are an artifact due to an electronic delay.
@@
To further characterize the nature of the strong nonlinear coupling, we measured the SHG at the single-photon power level by applying a weak signal at $\omega$ and monitoring the SHG signal at $2\omega$. The result is shown in Fig. 3. As expected, the SHG signal shown in Fig. 3a is more intense near the minimum of the avoided-level crossing and for flux offsets where the system is double resonant: $\tilde{\omega}_{3}=2\tilde{\omega}_{1}$, as confirmed (despite some shift due to some small discrepancy in the fitted energy levels) by the theoretical calculations in Fig. 3b (see Supplementary Information for details). The efficiency
$\eta\equiv S_{21}^{(2\omega)}/S_{21}^{(\omega)}$ (10)
of the SHG signal, shown in Fig. 3a, is about 0.1 at the point where the SHG amplitude is maximum. Such efficiency is achieved with an input corresponding to only 0.25 photons into the resonator, clearly operating in the quantum limit. To the best of our knowledge, SHG in the true (without additional many-photon drives) quantum limit has never been experimentally observed until now Yanagimoto2022Optica . A theoretical estimate (see Supplementary Fig. 3) shows that the efficiency can approach $\eta=0.3$ for 1.5 input photons. The efficiency can be significantly improved by reducing the internal losses of the resonator and setting the system at the perfect feeding condition so that no output signal at $\omega$ is present Bonizzoni2025 . Also, to the best of our knowledge, near-deterministic photon up/down conversion has so far been achieved predominantly using atomic level transitions Inomata2014PRL . Crucially, our approach has the potentiality to achieve near-deterministic pure photon up/down conversion without involving atomic transitions, thereby avoiding the associated losses.
Figure 3c shows that the SHG amplitude is thresholdless and starts with a linear dependence on the mean value of input photons $\bar{n}_{1}$, while it tends to saturate for higher input powers. It can clearly be distinguished from the noise floor even for the mean number of input photons in the resonator ${\overline{n}}_{1}$ significantly below $1$, which is a striking unprecedented result, made possible via the photonic strong coupling regime shown in Fig. 2b. The mean number of input photons in the resonator is evaluated using a standard input–output relationship Milburn2008Springer (see Supplementary Information for details). During the calibration of the photon number, the obtained fitted parameters allowed us to reproduce in very good agreement not only the data in Fig. 3c but also the spectra in Fig. 2 and the interference fringes in Fig. 4. Hence, although there might be a slight deviation from the true photon number, we expect it to be less than $10~\%$ of our fitted values. The hybridization between one- and two-photon states at the minimum of the avoided-level crossing gives rise to matrix elements ${X}_{+,1},\,{X}_{+,0},\,{X}_{1,0}$ all of the order of unity, which enable observation of the SHG well below the single-photon power level.
@@
A further direct consequence of the achieved photonic strong coupling is the observation of interference between a probe at $\omega$ ($2\omega$) and a control tone at $2\omega$ ($\omega$) both at the single-photon power level, resulting in the amplification or suppression of the transmitted amplitude of probe signals Astafiev2010PRL . The amplification and suppression are phase dependent, just as in the case of degenerate parametric amplifiers Tsai2008APL ; Lehnert2008NP . The difference is that a much stronger pump tone is not needed here, and we observe these interference phenomena at an incredibly low-power limit (average photon number in the resonator below one). The results are shown in Fig. 4. Changing the phase of the control tones, we can see that the transmitted amplitude of the signal at the $n=2$ ($n=1$) mode (signal field) is amplified or suppressed periodically with visibility approaching 1. It turns out that the visibility of the interference for the signal at $2\omega$ becomes relevant for higher intensities of the control signal (at $\omega$) (see Fig. 4c,d) with respect to what is observed for the signal at $\omega$ (see Fig. 4a,b). Conversely, to obtain a high visibility for interference at $\omega$, it requires a larger probe signal at $\omega$. This difference originates from the fact that the interference at $2\omega$ requires the stimulated conversion of two photons of the control tone. On the contrary, the probe signal at $\omega$ receives two photons from the down conversion of photons at $2\omega$ of the control tone. The observed tilting in the experimental data is attributed to a slight change in the phase of the signal when adjusting the output power. These two-tone interference data are well reproduced (except, of course, for the tilting) by numerical calculations based on the generalized master equations including a two-tone interaction term [see Supplementary Eq. (36)], shown in Fig. 4b,d.
We note that this work does not provide any experimental evidence of spontaneous down-conversion. This is a direct consequence of our setup, which selects only coherent signals, with the phase determined by the input signal. This technique allows to filter thermal noise and to allow the detection of very small signals. Spontaneous down conversion is not a coherent signal and it is thus filtered out. However, Fig. 4a provides clear evidence of stimulated down-conversion (phase-dependent amplification of the signal at $\omega$ in the presence of a tone at $2\omega$), which gives rise to a coherent signal that can be detected by our setup. A simple understanding of the phase sensitivity of the signal amplification at $\omega$ can be easily obtained considering the standard evolution equation for the linear parametric amplifier in the weak pump limit (see Supplementary Eq. 40).
@@ Discussion
We experimentally observed an unstimulated strong coupling between a single-photon and a two-photon Fock state. The observed single-photon strong nonlinear coupling is enabled by an embedded-in flux qubit, which, while remaining in its ground state, acts as an effective coupler between two modes of a superconducting coplanar waveguide resonator. The strong coupling between photon states of opposite parity is enabled by the spontaneous mixing of even- and odd-number photon states, which is a unique feature of the generalized quantum Rabi Hamiltonian in the ultrastrong coupling regime Nori2017PRA ; Wang2023NC .
The counterpart in time-resolved experiments of the quantum Rabi splitting reported here are quantum Rabi oscillations enabling the conversion of a single photon into two photons and vice versa with a period $T_{R}=2\pi/g_{\rm eff}$ and with an efficiency that can potentially approach $100\%$ Nori2017PRA (for $g_{\rm eff}$ significantly larger than the losses). In contrast to most conventional implementations of nonlinear optics, the regime demonstrated here can reach unit efficiency only using a minimal number of photons and without exciting atomic transitions, which introduce additional losses. For example, in many cases pure dephasing in superconducting artificial atoms is the main source of decoherence.
These results unlock a new regime for photons in quantum nonlinear optics Lukin2014NPh ; Bouwmeester2006PRL (coherent interaction between photon multiplets), which has important implications in both fundamental research Nori2017PRA and applications Zeilinger2011Nature ; Niu2018PRL . Examples of direct applications include
  1. (i)
    the realization of deterministic quantum light (single photon and photon pair) sources as in the cases of traditional spontaneous parametric down conversion (SPDC) but with an unprecedented efficiency (approaching $100\%$);
  2. (ii)
    an all-optical deterministic controlled-phase gate ($|\widetilde{0,1}\rangle\leftrightarrow-|\widetilde{0,1}\rangle$, a geometric phase effect under a full Rabi oscillation ($2T_{R}$), where the presence or the absence of a photon in mode 1 corresponds to the control qubit).
This two-photon quantum phase gate represents a key element for the realization of a still elusive universal nonlinear optical quantum computing (NOQC) Zeilinger2011Nature ; Niu2018PRL . These controlled quantum Rabi oscillations can be realized by using suitable flux-offset pulses to tune the system in and out the minimum of the avoided level crossing $|\widetilde{0,1}\rangle\leftrightarrow|\widetilde{2,0}\rangle$.
@@ Methods
Our device is a 3 mm $\times$ 10 mm chip fabricated on a single crystal sapphire substrate (see Supplementary Fig. 1b). The superconducting coplanar waveguide resonator is defined by patterning a niobium thin film using optical lithography. The flux qubit consists of two identical large JJs and a smaller JJ (with an area ratio $\alpha\approx\mathrm{0.5}$) in a loop geometry.
The qubit loop and its connection to the center conductor of the coplanar waveguide resonator are made by electron beam lithography and double-angle evaporation of aluminum. Placing the qubit loop near the antinode of the current distribution for the $n=2$ mode of the resonator makes the qubit coupled to the $n=1$ mode and the $n=2$ mode of the resonator simultaneously (see Fig. 1a).
The values of the Josephson energy $E_{J}/h=173\ \mathrm{GHz}$, the area ratio $\alpha\approx\mathrm{0.5}$ and ${E_{J}}/{E_{C}}\approx 100$ of the JJs in the qubit are optimized to ensure the flux qubit working as an effective nonlinear coupler, i.e., both far-away detuned and strongly coupled with the relevant modes of the resonator, and are used to numerically diagonalize the three-junction flux qubit Hamiltonian to estimate the values of $\Delta$ and $I_{p}$.
The device is coupled to the environment through two capacitors of different sizes (see Fig. 1). One of these input-output ports is used as an input port to excite the device with one or two coherent weak tones, while the other port is used to detect signals. In order to reduce the output through the input port, the size of the output capacitor has been chosen to be larger ($\simeq 1.7$ times) than that of the input one. Notice that the loss rate associated to a capacitive port is proportional to the square of the corresponding capacitance.
@@
The data that support the findings of this study are available at Code Ocean (https://doi.org/10.24433/CO.4816683.v1 CO ).
@@
The code that support the findings of this study are available at Code Ocean (https://doi.org/10.24433/CO.4816683.v1 CO ).
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China (Grant No. 2022YFA1405200), the National Natural Science Foundation of China (Grants No. 92265202, No. 92365210, No. 62074091, No. 12574547, No. U2230402, No. 12374325, and No. 12304387), and Tsinghua University Initiative Scientific Research Program. A.R. acknowledges the QuantERA grant SiUCs (Grant No. 731473), the PNRR MUR project PE0000023-NQSTI, and the ICSC - Centro Nazionale di Ricerca in High-Performance Computing, Big Data and Quantum Computing. Y.L. acknowledges the support of Beijing Municipal Science and Technology Commission (Grant No. Z221100002722011), and Young Elite Scientists Sponsorship Program by CAST (Grant No. 2023QNRC001). S.S. acknowledges the Army Research Office (ARO) (Grant No. W911NF1910065). F.N. is supported in part by: the Japan Science and Technology Agency (JST) [via the CREST Quantum Frontiers program Grant No. JPMJCR24I2, the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&D Grant No. JPMJMS2061].
Author contributions
S.-P.W., T.L. and J.Q.Y. conceived the experiment. Y.W., M.C. and W.W. fabricated the sample. S.-P.W. performed the experiment. S.-P.W., A.M., A.R., and S.S. analyzed and interpreted the data. A.M., A.R., F.N., and S.S. developed the theory. S.-P.W., A.M., A.R., Y.L., H.S., T.L., F.N., S.S. and J.Q.Y. discussed the results and contributed to the writing of the manuscript.
Competing Interests
The authors declare no competing interest.
Additional information
Correspondence and requests for materials should be addressed to Tiefu Li, Salvatore Savasta, or J. Q. You.